Recent Log Entries

Faraday Cup #1, ABIL tag #3789, Expected Removal Date changed.

Log Number: 3814807Logbook(s): Body: The following Bypassed Interlock had it's Expected Removal Date changed in the Accelerator Bypassed-Interlocks Log (ABIL) Previous Expected Removal Date: 22 Jul 2020 New Expected Removal Date: 30 Sep 2020 Who made the change: Terry Carlino Justification for the change: Repairs will be conducted during the SAD after the summer run.

CREW CHIEF SIGN IN

Log Number: 3814802Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 07-15-20
Time : 07:05:52
Day : Wednesday
Shift : Day

Crew Chief : Carlino
Program Deputy: McGuckin
Team : X
Arm : Carlino
Designated SSO: Carlino
LERF Operator:
Operators :

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

Box Supplies Were Turned Off

Log Number: 3814799Logbook(s): Tags: Body: 

The following box supplies were TURNED OFF by version 2-7 of the Box Supply Cycling Script:
ARC1A Ramp Rate = 8.113 A/s
ARC2A Ramp Rate = 7.176 A/s
ARC3A Ramp Rate = 10.228 A/s
ARC4A Ramp Rate = 11.047 A/s
ARC5A Ramp Rate = 10.408 A/s
ARC6A Ramp Rate = 9.652 A/s
ARC7A Ramp Rate = 10.471 A/s
ARC8A Ramp Rate = 11.950 A/s
ARC9A Ramp Rate = 9.456 A/s
ARCAA Ramp Rate = 12.604 A/s
DOG1E Ramp Rate = 3.831 A/s
DOG2E Ramp Rate N/A
DOG3E Ramp Rate N/A
DOG4E Ramp Rate N/A
DOG5E Ramp Rate N/A
DOG6E Ramp Rate N/A
DOG7E Ramp Rate N/A
DOG8E Ramp Rate N/A
DOG9E Ramp Rate N/A
LAM1C

1L12 shield not controlling well

Log Number: 3814796Logbook(s): Entry Makers: beneschBody: 

See bottom panel. JT and outlet temp are all over the place. 2K JT isn't as bad, but minimum of zero might be increased. And VIP1L1220LOG is troubling.

COMPLETED Install Hall C Turbo LCW Cooling Lines

Log Number: 3814795Logbook(s): Body: 

LCW hoses run and flow established. I would recommend
revisiting this work during SAD and installing a more robust
connection to the turbo. Presently it is a bard fitting and hose
clamps. Also installing isolation valves to all of the LCW taps
on the mezzanine.

COMMENT ON Crate Building

Log Number: 3814791Logbook(s): Body: 

Transform shipping crates requested by Rick Nelson, charge code
is RSRRF. Transformers presently located in NAB will need to be
moved to Bldg 36.

Back phasing bunch length measurement data

Log Number: 3814790Logbook(s): Entry Makers: kazimiBody: 

The bunch length measurement by back phasing in 4D00 spectrometer was done for Lasers A, B and C beams.
A save was done at the start of the work 21430 and another at the end 21433.
Fig 1: Initial rf setup
Fig 2,3: -90,+90 laser A
Fig 4,5: -91,+91 laser A
Fig 6: -90,+90 Harp laser A
Fig 7: -91, +91 Harp laser A
Fig 8: -90,-90 laser A
Fig 9: +90,+90 laser A
Fig 10: -90,-90 laser A Harp
Fig 11: +90,+90 laser A Harp
Fig 12: +90,+90 laser B Harp
Fig 13: +90,+90 laser C Harp
Fig 14: -90,-90 laser B Harp
Fig 15: -90,-90 laser C Harp
Fig 16: -90,-90 laser B
Fig 17: -90,-90 laser C

I will do the anal

Degaussing 4D00 magnet

Log Number: 3814789Logbook(s): Entry Makers: kazimiBody: 

I notice that correctors around 4D spectrometer have very high horizontal values. Degaussing the spectrometer magnet (Figure 1) and re-steering the line (Figure 2) lowered the corrector settings (Figure 3).
I know that our spectrometer magnet cycling routine for "Straight Ahead" needs some special dance to go the -4.8 Amps but I can not remember it exactly. I will email this to Gary Croke and Alicia. I would be good to fix the problem so that it works automatically.
I have left the magnet setting at setpoint #2 after degaussing (not the normal setpoint #1).

Autolog save

Log Number: 3814783Logbook(s): Tags: Body: saver ran from opsl632

Saved: kazimi 07-14-20 23:58:55 {A:1 B:5 C:5 E:1050 A=BS B=BS C=BS D=BS 0R08=0.00 0R08 dump in.

Autolog save

Log Number: 3814782Logbook(s): Tags: Body: saver ran from opsl632

Saved: kazimi 07-14-20 23:58:24 {A:1 B:5 C:5 E:1050 A=Tune B=BS C=BS D=BS 0R08=0.00 0R08 dump in.

CREW CHIEF SIGN IN

Log Number: 3814780Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 07-14-20
Time : 23:48:29
Day : Tuesday
Shift : Owl

Crew Chief : Frierson
Program Deputy: McGuckin
Team : D
Arm : RADCON
LERF Operator:
Operators :

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

MQD0R09 -- No Trim Card in Slot!

Log Number: 3814779Logbook(s): Entry Makers: moserBody: 

Linked log explains how MQD0R09 was not successfully cycling. Appeared calibration of card was off. Went to swap the card with a ready spare. Pulled the old trim card out and attempted to install a new one, but after over a dozen attempts could not get the card to properly seat with the backplane. The top rail appeared to be loose and wobbly about halfway into the rack.

Tried the original card again. First shot, went right in. Pulled it back out and compared the two --no obvious geometric differences. Tried the new card again, no luck.

COMPLETED ^LERF PSS Certification

Log Number: 3814778Logbook(s): Body: 

Completed Gun and Laser Chassis Functional Tests.

RF201, 202, 203 and Buncher were not tested.
This equipment will remain locked up until troubleshooting and repairs are complete and RF is needed.

W501B11 Negative Bulk Power Supply Tweaked

Log Number: 3814777Logbook(s): Entry Makers: moserBody: 

MQA{4, 6, 8}A21, off of W501B11 were failing to successfully cycle due to not making it to HMIN. Negative bulk power supply was outputting 28.4V, adjusted to 30.0. Quads cycling fine now.

MQA3A29 Trim Card Swapped

Log Number: 3814776Logbook(s): Entry Makers: moserBody: 

Magnet setpoint was stuck. Not taking demand, not cycling. Swapped trim card, all is well.

Pages